# Socially Aware Robot Navigation through Scoring Using Vision-Language Models

### v1

@@ Abstract
We propose VLM-Social-Nav, a novel Vision-Language Model (VLM) based navigation approach to compute a robot’s trajectory in human-centered environments. Our goal is to make real-time decisions on robot actions that are socially compliant with human expectations. We utilize a perception model to detect important social entities and prompt a VLM to generate guidance for socially compliant robot behavior. VLM-Social-Nav uses a VLM-based scoring module that computes a cost term that ensures socially appropriate and effective robot actions generated by the underlying planner. Our overall approach reduces reliance on large datasets (for training) and enhances adaptability in decision-making. In practice, it results in improved socially compliant navigation in human-shared environments. We demonstrate and evaluate our system in four different real-world social navigation scenarios with a Turtlebot robot. We observe at least $36.37\%$ improvement in average success rate and $20.00\%$ improvement in average collision rate in the four social navigation scenarios. The user study score shows that VLM-Social-Nav generates the most socially compliant navigation behavior.
@@ I Introduction
Mobile robots integrated into diverse indoor and outdoor human-centric environments are becoming increasingly prevalent. These robots serve various functions, ranging from package and food delivery [starship, scout] to service [dilligent] and home assistance [astro]. Overall, these roles necessitate interaction with humans and navigating seamlessly through public spaces with pedestrians. In such dynamic scenarios, it is important for the robots to engage in socially compliant interactions and navigation [xiao2022motion, mirsky2021conflict, mavrogiannis2023core].
@@
Challenges of social navigation: This paper focuses on the challenges of social navigation [mavrogiannis2023core] and addresses the ability of robots to navigate while adhering to social etiquette, especially contextual appropriateness, which requires robot to understand the relative importance of the context of the environment, tasks or interpersonal behaviors. However, navigating socially across varying contexts presents distinct challenges [francis2023principles, mavrogiannis2023core, mirsky2021conflict], including ensuring safety, comfort, politeness, and social norms.
@@
Inferring contextually appropriate navigation behaviors is challenging. Humans have various behaviors and the environmental or task contexts are also not easily to be categorized [mavrogiannis2023core]. A common strategy to handle the challenge is by learning-based approaches to empirically learn the complicated contexts. Imitation Learning (IL) is a recent emerging paradigm for desired navigation behavior [hirose2023sacson, xiao2020appld, xiao2022learning, panigrahi2023study, raj2024targeted]. This approach enables autonomous agents to navigate socially by learning from human demonstration. Other learning approaches, such as Reinforcement Learning (RL) have also been used to address this problem [kretzschmar2016socially]. While both methods demonstrate promising results in real-world settings, substantial datasets [thorDataset2019, karnan2022scand, nguyenmusohu] for training and reward engineering are required for their successful application.
@@
Language models are inherently well-suited for contextual understanding but not well applied in social navigation. Recent advances in Large Language Models (LLMs) and Vision-Language Models (VLMs) demonstrate a deep understanding of contextual information and have the potential to perform chain-of-thought [wei2022chain] and common sense reasoning [geva2021did, gpt4v2023]. Those processes are inherent to social navigation, especially the challenges of contextual appropriateness and politeness, which require understanding the task/environmental context and the behavior of humans. This capability has also been evaluated across diverse domains of robotics, including human-like driving scenarios [wen2023road, sha2023languagempc] and autonomous robot navigation [10160969, pmlr-v205-shah23b]. However, using language models for social navigation is not well explored and the language models suffer from high latency for real-time navigation, and the issue impedes the smoothness and efficiency of human-robot social interaction.
@@
Main Results: In this paper, we present VLM-Social-Nav, a new approach that uses VLMs to interpret contextual information from the robot observation to help autonomous agents improve their navigation abilities in human-centered environments. In particular, we leverage a VLM to analyze and reason about the current social interaction, and generate an immediate preferred robot action to guide the motion planner. Our VLM-based scoring module computes the social cost, which is used for the bottom-level motion planner to output the appropriate robot action. To overcome the limitation of existing VLMs’ latency issue, we utilize a state-of-art perception model (i.e., YOLO [YOLO2023]) to detect key entities that are used for social interactions (e.g., humans, gestures, and doors), query a VLM to generate socially compliant navigation behavior, and compute the social cost. We demonstrate VLM-Social-Nav in four different indoor scenarios with human interactions. Unlike previous social navigation approaches, VLM-Social-Nav can better deal with social navigation scenarios by interpreting the situation based on common sense, through observation without any training with a large dataset. Some of our main results include:
  • We propose VLM-Social-Nav, a novel approach for social robot navigation by integrating VLMs with optimization-based or scoring-based local motion planners and a state-of-the-art perception model. This enables robots to detect social entities efficiently and make real-time decisions on socially compliant robot behavior.
  • We propose a VLM-based scoring module that translates the current robot observation and the textual instructions into a relevant social cost term. This cost term is used for the bottom-level motion planner to output appropriate robot action.
  • We evaluate VLM-Social-Nav in four different real-world indoor social navigation scenarios. We measure the success rate and the collision rate. We also perform a user study and compare the results with a Dynamic-Window Approach (DWA) [fox1997dynamic] and Behavior Cloning (BC) [pomerleau1988alvinn] method trained on a state-of-the-art large Socially CompliAnt Navigation Dataset (SCAND) [karnan2022scand]. We achieve at least $36.37\%$ improvement in average success rate and $20.00\%$ improvement in average collision rate in the four social navigation scenarios. The user study score shows that VLM-Social-Nav generates the most socially compliant navigation behavior.
@@ II Related Work
In this section, we give an overview of existing works related to social navigation, including safety requirements of social navigation, different challenges of contextual appropriateness, and Large Foundation Models (LFMs) for robot navigation.
@@ II-A Safety Requirement of Social Navigation
For social navigation, keeping safety is the basic requirement for interaction with humans and navigation in dynamic challenging scenarios [van2010optimal, fox1997dynamic, liang2021vo, liang2021crowd]. DWA [fox1997dynamic] calculates the collision constraints of the robots’ actions and chooses the best feasible action closest to the target as the output. Model Predictive Control approaches [10341869, williams2017information] are also widely used for collision avoidance tasks, which provide smooth trajectories for the robot to follow; the learning-based Model Predictive Control (MPC) [williams2017information, xiao2022learning] enhances the performance of the original MPC method by empirically learning the best actions for navigation, which requires a large expert demonstration dataset for training. Velocity-Obstacle (VO)-based approaches are more efficient and can be used to simulate the actions of crowds [van2010optimal, best2014densesense], but these approaches don’t take into account the uncertainties in the perception output. PRVO [gopalakrishnan2017prvo] and OFVO [liang2021vo] handle the uncertainties of perception in motion planning, but those approaches require a hard threshold to set the confidence for planning. To deal with this issue, learning-based methods empirically train the policies by demonstrations [liang2021crowd, aradi2020survey, sun2021motion], whereas other methods [liang2021crowd, aradi2020survey] use reinforcement learning to train the robot in a simulator and implement it in real-world scenarios. However, learning-based approaches require a significant amount of data or on-policy training with realistic simulators to learn the task.
@@ II-B Contextual Appropriateness of Social Navigation
Researchers have proposed diverse methodologies to integrate social awareness into mobile robot navigation systems. The development of a comprehensive social navigation system poses inherent complexity, requiring sophisticated perception and reasoning capabilities to navigate environments shared with humans and other robots [mirsky2021conflict]. Establishing an accurate definition of social navigation is paramount, given its potential variation across cultures and platforms. Except for the safety requirement, assessing social compliance presents an additional challenge, contingent on the scenario and platform, and should adhere to Contextual Appropriateness, which includes Comfort, Legibility, Social Competency, Politeness, Understanding Other Agents, and Proactivity [gao2022evaluation, francis2023principles]. Various methodologies are employed to address this challenge, with a significant focus on enhancing learning methods through reinforcement learning, learning from demonstration—particularly by analyzing examples of human trajectories or robots operated by humans—and the utilization of simulated datasets [9679193, 5354147, 7539621, li2017role, nazeri2024vanp]. Additionally, various datasets have been collected for this purpose. Socially CompliAnt Navigation Dataset (SCAND) [karnan2022scand] and Multi-Modal Social Human Navigation Dataset (MuSoHu) [nguyenmusohu] are two recent large-scale social human navigation datasets in many natural human-inhabited public spaces for robots to learn similar, human-like, socially compliant navigation behaviors. Despite extensive research employing diverse approaches with machine learning techniques [10323465], less emphasis has been placed on formulating social navigation using VLMs, which successfully analyze the contextual information of the environment.
Figure 2: The overall system architecture of VLM-Social-Nav. Our real-world perception (vision) model detects important social entities (e.g., humans, gestures, and doors) in real time and prompts the VLM-based scoring module, indicated as in green parts, to compute social cost $C_{social}$, which is used to generate socially compliant robot action.
@@ II-C Large Foundation Models for Navigation
Recent advancements in Language Foundation Models (LFMs) [bommasani2022opportunities], encompassing Vision-Language Models (VLMs) and Language-Language Models (LLMs), show significant potential in robotic navigation, despite the challenges like bias and reliability [payandeh2023susceptible] associated with the LFMs. SayCan [Ahn2022] integrate LLMs for high-level task planning. GPT-Driver [mao2023gptdriver] evaluates the performance of GPT-3.5 in simulation for autonomous driving, framing motion planning as a language modeling problem. LL3MVN [Yu_2023] constructs semantic maps of environments and utilizes LLMs to reach long-term goals, while LLaDA [li2024driving] enables autonomous vehicles to adapt to diverse traffic rules across regions. LM-Nav [shah2022lmnav] utilizes GPT-3 and CLIP [radford2021learning] to navigate outdoor environments based on natural language instructions, combining language and visual cues for optimal path planning. RT-1 [brohan2022rt1] and RT-2 [brohan2023rt2] models leverage multimodal robotics data for navigating complex environments and offer greater generalizability through larger dataset training [shah2022lmnav, xu2023vision, brohan2022rt1, brohan2023rt2]. However, language models are not well explored for social navigation, even though VLMs and LLMs are powerful for understanding contextual information. We bridge the gap and propose a novel approach for social navigation by taking advantage of the contextual understanding capability of VLMs.
@@ III-A Problem Definition
Navigation is the task of generating and following an efficient collision-free path from an initial location to a goal [mavrogiannis2023core]. In general, the overall system consists of a global planner and a local planner. A global planner is designed to find a collision-free path to reach a goal, while a local planner aims to navigate the robot through its immediate surroundings, making real-time adjustments to deal with vehicle dynamics and around obstacles.
For social robot navigation, humans are no longer perceived only as dynamic obstacles but also as social entities [rios2015proxemics, mirsky2021conflict]. It necessitates integrating social norms into robot behaviors. We define the social robot navigation problem as a Markov Decision Process (MDP): $\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{C}\rangle$ where $\mathbf{s}=(x,y,\theta)\in\mathcal{S}$ is a state consisting of a robot pose, $\mathbf{a}=(v,w)\in\mathcal{A}$ is an action consisting of a linear and an angular velocity of a robot, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function characterizing the dynamics of the robot, and $\mathcal{C}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is a cost function. Given a cost function $\mathcal{C}$, the motion planner finds $(v^{*},w^{*})$ that minimizes the expected cost. The cost function takes the following form:
$C(\mathbf{s},\mathbf{a})=\alpha\cdot C_{goal}+\beta\cdot C_{obst}+\gamma\cdot C_{social},$ (1)
where $C_{goal}$ encourages movement toward the goal, $C_{obst}$ discourages collisions with obstacles, and $C_{social}$ encourages the robot to follow the social norms. $\alpha$, $\beta$, and $\gamma$ are a non-negative weight for each cost term.
The social cost term $C_{social}$ encompasses various factors that govern human-robot interactions in shared environments. Defining them mathematically poses challenges. For VLM-Social-Nav, we define $C_{social}$ as:
$C_{social}=\|\mathcal{B}-\mathcal{B}_{h}\|,$ (2)
where $\mathcal{B}$ is a navigation behavior, and $\mathcal{B}_{h}$ is a navigation behavior humans would adopt in accordance with social conventions. Minimizing the deviation between them will encourage the robot to emulate socially acceptable human behavior. While $\mathcal{B}_{h}$ can be obtained through various methods, including large datasets [thorDataset2019, karnan2022scand, nguyenmusohu], we leverage the power of VLM to compute appropriate behavior based on a rich contextual understanding and nuanced interpretations from perceived images and given prompts. We elaborate further in Section III-C.
@@ III-B VLM-based Social Navigation Architecture
Figure 2 highlights the overview of VLM-Social-Nav. Our formulation is built upon an autonomous navigation system comprising a global path planner and a local motion planner. Our local motion planner considers the sensor inputs and outputs a robot action that minimizes the cost function $C$.
To improve navigation abilities in a social-interaction context, we introduce a VLM-based scoring module. As the robot navigates through the environment, our real-world perception module detects social entities, such as humans, gestures, and doors, in real time. When important social cues are detected, our VLM-based scoring module is prompted and outputs a socially acceptable robot action based on the current robot observation. It is calculated as a cost term to be used by the motion planner. VLM-Social-Nav is designed to make real-time decisions on robot actions based on the current observations. The social cost term is integrated only when it is necessary, i.e., when there is any human interaction involved.

### v2

# VLM-Social-Nav: Socially Aware Robot Navigation through Scoring using Vision-Language Models

@@ Abstract
We propose VLM-Social-Nav, a novel Vision-Language Model (VLM) based navigation approach to compute a robot’s motion in human-centered environments. Our goal is to make real-time decisions on robot actions that are socially compliant with human expectations. We utilize a perception model to detect important social entities and prompt a VLM to generate guidance for socially compliant robot behavior. VLM-Social-Nav uses a VLM-based scoring module that computes a cost term that ensures socially appropriate and effective robot actions generated by the underlying planner. Our overall approach reduces reliance on large training datasets and enhances adaptability in decision-making. In practice, it results in improved socially compliant navigation in human-shared environments. We demonstrate and evaluate our system in four different real-world social navigation scenarios with a Turtlebot robot. We observe at least $27.38\%$ improvement in the average success rate and $19.05\%$ improvement in the average collision rate in the four social navigation scenarios. Our user study score shows that VLM-Social-Nav generates the most socially compliant navigation behavior.
Index Terms:
Motion and Path Planning, Task and Motion Planning, Integrated Planning and Control
@@ I Introduction
Mobile robots integrated into diverse indoor and outdoor human-centric environments are becoming increasingly prevalent. These robots serve various functions, ranging from package and food delivery [starship, scout] to service [dilligent] and home assistance [astro]. Overall, these roles necessitate interaction with humans and navigating seamlessly through public spaces with pedestrians. In such dynamic scenarios, it is important for the robots to engage in socially compliant interactions and navigation [xiao2022motion, mirsky2024conflict, mavrogiannis2023core].
@@
This paper focuses on the challenges of social navigation [mavrogiannis2023core]. It addresses the ability of robots to navigate while adhering to social etiquette, especially contextual appropriateness, which requires robots to understand environment contexts, current tasks, and interpersonal behaviors. Therefore, navigating socially across varying contexts presents distinct challenges [francis2023principles, mavrogiannis2023core, mirsky2024conflict], including ensuring safety, comfort, and politeness, as well as adhering to social norms.
@@
Inferring contextually appropriate navigation behaviors is challenging. Humans have various behaviors and the environmental or task contexts also cannot be easily categorized [mavrogiannis2023core]. A common strategy to handle the challenge is by learning-based approaches to empirically learn the complicated contexts. Imitation Learning (IL) is a recent emerging paradigm for desired navigation behavior [hirose2023sacson, xiao2020appld, xiao2022learning, panigrahi2023study, raj2024targeted]. This approach enables autonomous robots to navigate socially by learning from human demonstrations. Other learning approaches, such as Reinforcement Learning (RL) have also been used to address this problem [kretzschmar2016socially]. While both methods demonstrate promising results in real-world settings, substantial datasets [thorDataset2019, karnan2022scand, nguyenmusohu] for training and reward engineering are required for their successful application and it is hard to generalize.
@@
Language models are inherently well-suited for contextual understanding but not well applied in social navigation. Recent Large Language Models (LLMs) and Vision-Language Models (VLMs) demonstrate a deep understanding of contextual information and have the potential to perform chain-of-thought [wei2022chain] and common sense reasoning [geva2021did, gpt4v2023]. Those processes are inherent to social navigation, especially the challenges of contextual appropriateness and politeness, which require understanding the task/environmental context and the behavior of humans. This capability has also been evaluated across diverse domains of robotics, including human-like driving scenarios [wen2024road, sha2023languagempc] and autonomous robot navigation [10160969, shah2022lmnav]. However, using language models for social navigation is not well explored and the language models suffer from high latency for real-time navigation, and the issue impedes the smoothness and efficiency of human-robot social interaction.
@@
Main Results: In this paper, we present VLM-Social-Nav, a new approach that uses VLMs to interpret contextual information from robot observation to help autonomous robots improve their navigation abilities in human-centered environments. In particular, we leverage a VLM to analyze and reason about the current social interaction, and generate an immediate preferred robot action to guide an underlying motion planner. Our VLM-based scoring module computes a social cost, which is used for the bottom-level motion planner to output the appropriate robot action. To overcome the limitation of existing VLMs’ latency issue, we utilize a state-of-the-art perception model (i.e., YOLO [YOLO2023]) to detect key entities that are used for social interactions (e.g., humans, gestures, and doors) and query a VLM to generate socially compliant navigation behavior and compute the social cost. We demonstrate VLM-Social-Nav in four different indoor scenarios with human interactions. Unlike previous social navigation approaches, VLM-Social-Nav can better navigate through social scenarios by interpreting the situation based on common sense without any dedicated training on a large dataset. Some of our main results include:
  • We propose VLM-Social-Nav, a novel approach for social robot navigation by integrating VLMs with optimization-based or scoring-based motion planners and a state-of-the-art perception model for better VLM efficiency.
  • We propose a VLM-based scoring module that translates the current robot observation and textual instructions into a relevant social cost term. This cost term is used for the bottom-level motion planner to output appropriate robot action.
  • We evaluate VLM-Social-Nav in four different real-world indoor social navigation scenarios along with a user study and compare the results with a Dynamic-Window Approach (DWA) [fox1997dynamic] and Behavior Cloning (BC) [pomerleau1988alvinn] method trained on a state-of-the-art large Socially CompliAnt Navigation Dataset (SCAND) [karnan2022scand]. VLM-Social-Nav achieves at least $27.38\%$ improvement in average success rate and $19.05\%$ improvement in average collision rate in the four social navigation scenarios. The user study score shows that VLM-Social-Nav generates the most socially compliant navigation behavior.
@@ II Related Work
In this section, we give an overview of existing works related to safety requirements and different challenges of contextual appropriateness in social robot navigation, and Large Foundation Models (LFMs) for robot navigation.
@@ II-A Safety Requirement of Social Navigation
For social navigation, safety is a basic requirement for interaction with humans and navigation in dynamic challenging scenarios [van2010optimal, fox1997dynamic, liang2021vo, liang2021crowd]. DWA [fox1997dynamic] calculates the collision constraints of the robots’ actions and chooses the best feasible action closest to the target as the output. Model Predictive Control (MPC) approaches [10341869, williams2017information] are also widely used for collision avoidance tasks, which provide smooth trajectories for the robot to follow; learning-based MPC [williams2017information, xiao2022learning] enhances the performance of the original MPC method by empirically learning the best actions for navigation, which requires a large expert demonstration dataset for training. Velocity-Obstacle (VO)-based approaches are more efficient and can be used to simulate the actions of crowds [van2010optimal, best2014densesense], but these approaches don’t take into account the uncertainties in the perception output. PRVO [gopalakrishnan2017prvo] and OFVO [liang2021vo] handle the uncertainties of perception in motion planning, but those approaches require a hard threshold to set the confidence for planning. To deal with this issue, learning-based methods empirically train the policies by demonstrations [liang2021crowd, aradi2020survey, sun2021motion], whereas other methods [liang2021crowd, aradi2020survey] use reinforcement learning to train the robot in a simulator and implement it in real-world scenarios. However, learning-based approaches require a significant amount of data or on-policy training with realistic simulators to learn the task. Social interactions are highly nuanced and context-dependent. Simulating these interactions accurately requires sophisticated models of human behavior and interaction dynamics, which is not trivial.
@@ II-B Contextual Appropriateness of Social Navigation
Researchers have proposed diverse methodologies to integrate social awareness into mobile robot navigation systems. The development of a comprehensive social navigation system poses inherent complexity, requiring sophisticated perception and reasoning capabilities to navigate environments shared with humans and other robots [mirsky2024conflict]. Establishing an accurate definition of social navigation is paramount, given its potential variation across cultures and platforms. Except for the safety requirement, assessing social compliance presents an additional challenge, contingent on the scenario and platform, and should consider contextual appropriateness [gao2022evaluation, francis2023principles]. Various methodologies are employed to address this challenge, with a significant focus on enhancing learning methods through reinforcement learning, learning from demonstration—particularly by analyzing examples of human trajectories or robots operated by humans—and the utilization of simulated datasets [9679193, 5354147, 7539621, li2017role, nazeri2024vanp]. Additionally, various datasets have been collected for this purpose. Socially CompliAnt Navigation Dataset (SCAND) [karnan2022scand] and Multi-Modal Social Human Navigation Dataset (MuSoHu) [nguyenmusohu] are two recent large-scale social human navigation datasets in many natural human-inhabited public spaces for robots to learn similar, human-like, socially compliant navigation behaviors. Despite extensive research employing diverse approaches with machine learning techniques [10323465], VLMs have not been used to solve the social navigation problem, which, however, have shown the potential to successfully analyze the contextual information of the environment.
Figure 2: The overall system architecture of VLM-Social-Nav. Our real-world perception (vision) model detects important social entities (e.g., humans, gestures, and doors) in real time and prompts the VLM-based scoring module to compute social cost $\mathcal{C}_{\textrm{social}}$, which is used to generate socially compliant robot action.
@@ II-C Large Foundation Models for Navigation
Recent advancements in Language Foundation Models (LFMs) [bommasani2022opportunities], encompassing Vision-Language Models (VLMs) and Large-Language Models (LLMs), show significant potential in robotic navigation, despite the challenges like bias and reliability [payandeh2023susceptible] associated with the LFMs. SayCan [Ahn2022] integrates LLMs for high-level task planning. GPT-Driver [mao2023gptdriver] evaluates the performance of GPT-3.5 in simulation for autonomous driving, framing motion planning as a language modeling problem. L3MVN [yu2023l3mvn] constructs semantic maps of environments and utilizes LLMs to reach long-term goals, while LLaDA [li2024driving] enables autonomous vehicles to adapt to diverse traffic rules across regions. LM-Nav [shah2022lmnav] utilizes GPT-3 and CLIP [radford2021learning] to navigate outdoor environments based on natural language instructions, combining language and visual cues for optimal path planning. Despite their powerful capabilities in contextual understanding and commonsense reasoning, language models have not been extensively investigated for social navigation. Our approach proposes a novel method to navigate robots in a socially compliant manner.
@@ III-A Problem Definition
Navigation is the task of generating and following an efficient collision-free path from an initial location to a goal [mavrogiannis2023core]. In general, the overall system consists of a global planner and a local planner. A global planner is designed to find a collision-free path to reach a goal, while a local planner aims to navigate the robot through its immediate surroundings, making real-time adjustments to deal with vehicle dynamics and surrounding obstacles.
For social robot navigation, humans are no longer perceived only as dynamic obstacles but also as social entities [rios2015proxemics, mirsky2024conflict]. It necessitates integrating social norms into robot behaviors. We define the social robot navigation problem as a Markov Decision Process (MDP): $\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{C}\rangle$ where $\mathbf{s}=(x,y,\theta)\in\mathcal{S}$ is a state consisting of a robot pose, $\mathbf{a}=(v,w)\in\mathcal{A}$ is an action consisting of a linear and an angular velocity of the robot, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function characterizing the dynamics of the robot, and $\mathcal{C}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is a cost function. Given a cost function $\mathcal{C}$, the motion planner finds $(v^{*},w^{*})$ that minimizes the expected cost. The cost function takes the following form:
$\mathcal{C}(\mathbf{s},\mathbf{a})=\alpha\cdot\mathcal{C}_{\textrm{goal}}+\beta\cdot\mathcal{C}_{\textrm{obst}}+\gamma\cdot\mathcal{C}_{\textrm{social}},$ (1)
where $\mathcal{C}_{\textrm{goal}}$ encourages movement toward the goal, $\mathcal{C}_{\textrm{obst}}$ discourages collisions with obstacles, and $\mathcal{C}_{\textrm{social}}$ encourages the robot to follow the social norms. $\alpha$, $\beta$, and $\gamma$ are a non-negative weight for each cost term.
The social cost term $\mathcal{C}_{\textrm{social}}$ encompasses various factors that govern human-robot interactions in shared environments. Defining them mathematically poses challenges. For VLM-Social-Nav, we define $\mathcal{C}_{\textrm{social}}$ as:
$\mathcal{C}_{\textrm{social}}=\|\mathcal{B}-\mathcal{B}_{h}\|,$ (2)
where $\mathcal{B}$ is a navigation behavior, and $\mathcal{B}_{h}$ is a navigation behavior humans would adopt in accordance with social conventions. Minimizing the deviation between them will encourage the robot to emulate socially acceptable human behaviors. While $\mathcal{B}_{h}$ can be obtained through various methods, including large datasets [thorDataset2019, karnan2022scand, nguyenmusohu], we leverage the power of a VLM to compute appropriate behavior based on its rich contextual understanding and nuanced interpretations from perceived images and given prompts. We elaborate further in Section III-C.
@@ III-B VLM-based Social Navigation Architecture
Figure 2 highlights the overview of VLM-Social-Nav. Our formulation is built upon an autonomous navigation system comprising a perception layer and an optimization-based motion planner. The motion planner considers the sensor inputs and outputs a robot action that minimizes the cost function $C$.
While LiDAR detects geometric information useful for obstacle avoidance, RGB images provide contextual details of the current environment. They contain rich information crucial for social navigation. To enhance navigation capabilities within social contexts, we propose a VLM-based scoring module. VLMs excel in contextual understanding, interpreting scenes not solely based on visual features but also considering social dynamics. VLMs generate socially appropriate robot actions based on current observations and input instructions. Our VLM-based scoring module then calculates a cost term to be used by the motion planner.
While VLMs can generate navigation behaviors that comply with social norms, continuously querying large VLMs for new responses is prohibitively computationally expensive for real-time navigation. To address this challenge, we incorporate a real-time perception model. This model identifies social entities such as humans, gestures, and doors as the robot navigates its environment. Our VLM-based scoring module activates only when significant social cues are detected, ensuring that the social cost term is integrated only when necessary, i.e., when there is any human interaction involved. This approach reduces the VLM queries and facilitates real-time navigation efficiency for our approach. Algorithm 1 summarizes an overview of our VLM-Social-Nav process.
Input : RGB image $\mathcal{I}$, LiDAR point cloud $\mathcal{L}$, prompt $\mathcal{P}$, goal position $\mathbf{p}_{g}$
1 Initialize robot position $\mathbf{p}_{r}$;
2 while not at goal position $\mathbf{p}_{r}\neq\mathbf{p}_{g}$ do
3       $\mathcal{I}\leftarrow$ Read image sensor data;
4       $\mathcal{L}\leftarrow$ Read LiDAR sensor data;
5       $e\leftarrow\text{Perception Model}(\mathcal{I})$;
6       $(\delta_{s},\delta_{d})\leftarrow 0$;
7       if social entities detected in $e$ then
8             $\mathcal{B}_{h}=\text{VLM}(\mathcal{I},\mathcal{P},\mathbf{a})$;
9             $\mathcal{B}_{h}\mapsto(\delta_{s},\delta_{d})$;
10
11       forall possible actions $\mathbf{a}$ do
12             $\mathcal{C}_{\textrm{social}}=\text{VLM-based scoring}(\delta_{s},\delta_{d})$;
13             Calculate the total cost $\mathcal{C}=\alpha\cdot\mathcal{C}_{\textrm{goal}}+\beta\cdot\mathcal{C}_{\textrm{obst}}+\gamma\cdot\mathcal{C}_{\textrm{social}}$;
14
15       end forall
16      Find action $\mathbf{a}$ with minimal cost $\mathcal{C}$ and execute;
17       Update robot position $\mathbf{p}_{r}$;
18
19 end while
Algorithm 1 VLM-Social-Nav